OpenStudy (anonymous):

How to work on this, 2(2x^2 + y^2)dx - xydy =0

OpenStudy (primeralph):

http://tutorial.math.lamar.edu/Classes/DE/DE.aspx

OpenStudy (anonymous):

this is a toughy

OpenStudy (tkhunny):

\(M(x,y) = 2(2x^2 + y^2)\) \(N(x,y) = -xy\) \(M_{y}(x,y) = 4y\) \(N_{x}(x,y) = -y\) That's not quite exact, but it is a clue. \(\dfrac{M_{y} - N_{x}}{N(x,y)} = -5/x = Q(x)\) \(\int Q(x)\;dx = -5\cdot \ln|x| + C\) \(e^{-5\cdot \ln|x| + C} = \dfrac{A}{x^{5}} = u(x)\) Now, see if it's exact. \(M(x,y)\cdot u(x) = \) \(N(x,y)\cdot u(x) = \)